OpenStudy (anonymous):

\[\int\limits_{1}^{6} x^2/\sqrt{x+3}\]

OpenStudy (anonymous):

how do i evaluate that integral

OpenStudy (raden):

int by subs, let u=x+3

OpenStudy (unklerhaukus):

\[\int_1^6\frac{x^2}{\sqrt{x+3}}\text dx\] \(u=x+3\) \(\frac{\text du}{\text dx}=1\) \(\text du=\text dx\) \[=\int_1^6\frac{(u-3)^2}{\sqrt{u}}\text du\]

OpenStudy (unklerhaukus):

\[=\int_1^6\frac{u^2-6u+9}{\sqrt{u}}\text du\] \[=\int_1^6\frac{u^2}{\sqrt{u}}\text du-6\int_1^6\frac{u}{\sqrt{u}}\text du+9\int_1^6\frac{1}{\sqrt{u}}\text du\]

OpenStudy (lgbasallote):

^?

OpenStudy (unklerhaukus):

no?